# Asymptotic stability equals exponential stability—while you twist your eyes

### v1

@@ Abstract
Keywords—dynamical systems, homotopy, stability
@@ 1   Introduction
In the context of what we call today Conley index theory (see Section A), Charles C. Conley himself posed the following “converse question” in the late 1970s: “To what extent does the homotopy index [Conley index] itself determine the equivalence class of isolated invariant sets which are related by continuation?” [ref:conley1978isolated, p. 83]. Then, recently, Matthew D. Kvalheim proved that uniquely integrable $C^{0}$ vector fields, on a $C^{\infty}$ manifold $M$, rendering a compact set $A\subseteq M$ asymptotically stable, are homotopic on an open neighbourhood $U\supseteq A$ such that throughout the homotopy the vector fields do not vanish on $U\setminus A$ [ref:kvalheim2022obstructions, Thm. 1]. Connecting this result to Conley’s question, a follow-up question (revitalized) by Kvalheim—which is the central question of this note—is the following:
Question:Are, dynamical systems that render a set $A$ asymptotically stable, homotopic through dynamical systems that preserve this notion of stability?
This question, in one form or another, inspired several works, for instance, [ref:reineck1992continuation, ref:mrozek2000conley, ref:JongeneelSchwan2024TAC]. Here, we will elaborate on commentary by the author in [ref:JongeneelSchwan2024TAC]. Specifically, in the seminal paper “Asymptotic stability equals exponential stability, and ISS equals finite energy gain—if you twist your eyes” from 1999, Lars Grüne, Eduardo D. Sontag and Fabian R. Wirth showed that asymptotic stability “equals” exponential stability in the sense that if an equilibrium point is asymptotically stable under some vector field on $\mathbb{R}^{n}$, then, there is a suitable change of coordinates rendering this point exponentially stable [ref:grune1999asymptotic]. Such a change of coordinates is understood to be instantaneous. However, by leveraging their work, we show in this note that asymptotic stability can (almost) always be continuously “transformed” into exponential stability, while preserving asymptotic stability throughout the transformation, see Theorem 3.4. Differently put, asymptotic stability equals exponential stability—not only if you twist your eyes, but while you twist your eyes. The key to this is to observe that we can in fact select the transformation of Grüne, Sontag and Wirth to be an element of the orientation-preserving homeomorphism group on $\mathbb{R}^{n}$, not just any homeomorphism cf[ref:JongeneelSchwan2024TAC, Sec. III], see the proof of Proposition 3.2.
This result provides a partial solution to Conley’s converse question as it turns out that the asymptotically stable systems under consideration can be continuously transformed into the same exponentially stable system and hence, by transitivity, into each other. Concurrently, we emphasize in this note that results of this form typically extend to discontinuous vector fields when solutions are understood in the sense of Filippov. We also discuss intimate connections with optimization (e.g., see Example 3.2) and optimal transport (e.g., see Example 4.1).
@@ 1.1   Related work
It can be argued that questions of the form above emerged from studies aimed at classifying manifolds, maps, vector fields and so forth. A successful, yet coarse, resolution has been found in the study of these objects up to homotopy, e.g., motivated early on by the fundamental group being homotopy invariant, Hopf’s degree theorem, CW complexes, intractability of topological equivalence and more work at the intersection of topology and dynamical systems.
We cannot do justice here to the wealth of work in this area, but let us mention that inspired by Aleksandr Andronov, Lev Pontryagin, René Thom, Maurico Peixoto, George D. Birkhoff, John Milnor and several others, it was in particular Stephen Smale highlighting that this intersection is an interesting one for both topology and dynamical systems. As put by Sheldon Newhouse: “… one of the great influences that Steve had, at least on me and I think many others, aside from excitement, was the whole idea that one might get a structure theory for general dynamical systems. That such a general structure theory for most systems could exist I think was a goal of Poincare and Birkhoff, and somehow in the intervening years, it was lost sight of by people working on specialized problems. The idea of getting a global structure of all systems has spawned a lot of development in the past, and it is still going on.” [ref:hirsch2012topology, p. 183].
@@
Indeed, after seminal work by Smale on the qualitative classification of dynamical systems [ref:smale1967differentiable], several homotopy results appeared being concerned with Morse-Smale vector fields, e.g., see [ref:asimov1975homotopy, ref:newhouse1976, ref:franks1979morse]. Similarly, one can study if gradient vector fields are homotopic through gradient vector fields, e.g., while restricting equilibria [ref:parusinski1990gradient], see also [ref:reineck1991continuation, ref:kvalheim2022obstructions] for work close in spirit to ours. These works were typically concerned with structural stability and the study of the topology of spaces of dynamical systems, that is, to understand when dynamical systems are in some sense close or equivalent, e.g., see [ref:shub1975homology] for pointers and further results.
@@
Then, in control theory, these tools (i.e., certain homotopy invariants) were used to construct necessary conditions for feedback controllers to exist, e.g., if a desirable (closed-loop) dynamical system belongs to a certain homotopy class, then, there must be at least exist some feedback that renders the control system a member of this class, regardless of it resulting in the actual desirable system, e.g., see [ref:brockett1983asymptotic, ref:krasnosel1984geometrical, ref:zabczyk1989, ref:Coron1990].
@@
Most of these results have in common that the original objects are homotoped to something simple, something “canonical” where we do our analysis and computations. For instance, in the context of control systems, when our goal is stabilization of the origin on $\mathbb{R}^{n}$ through feedback, then, the canonical differential equation is $\dot{x}=-x$ and we would like our closed-loop dynamical system to be in some qualitative sense equivalent to this equation. This is precisely how the “index condition” by Krasnosel’skiĭ and Zabreĭko is derived [ref:krasnosel1984geometrical, Sec. 52]. Clearly, necessary conditions of this form are only as valuable as the homotopy class is distinctive, e.g., although the index condition is powerful, it cannot differentiate between $\dot{x}=x$ and $\dot{x}=-x$ on $\mathbb{R}^{2}$. This is precisely what motivates us, we expect that understanding Conley’s converse question can lead to stronger necessary conditions for continuous stabilizing feedback to exist.
@@
Besides possibly stronger necessary conditions for continuous stabilization, the study of Conley’s converse question directly relates to understanding topological properties of spaces of stable dynamical systems (as was partial motivation for our previous work [ref:JongeneelSchwan2024TAC]). This is of practical importance as several frameworks in optimal control and reinforcement learning aim to optimize over precisely such a space, e.g., moving from a pre-stabilized system to an optimally controlled system, we point to [ref:berkenkamp2017safe, ref:Fazel_18, ref:wang2022learning, ref:furieri2024learning] and references therein.
@@
Before detailing further technicalities, we recall that Conley’s converse question is not trivial.
@@
Both $A(0)$ and $A(1)$ correspond to global asymptotically stable systems, yet, for $s=(1/2)$ we find that the system $\dot{x}=A(s)x$ is unstable. Hence, we cannot just construct straight-line homotopies between stable vector fields and expect that stability is preserved. We know from [ref:JongeneelSchwan2024TAC] that instead, for linear vector fields we should homotope via the canonical ODE $\dot{x}=-x$. Explicitly, one could consider the following path of linear vector fields defined by
$H(x;s):=\begin{cases}\begin{pmatrix}-1&(1-2s)\cdot 10\\ 0&-1\end{pmatrix}x\quad&s\in[0,1/2]\\ \begin{pmatrix}-1&0\\ (2s-1)\cdot 10&-1\end{pmatrix}x\quad&s\in(1/2,1]\end{cases}.$
$\circ$
Notation and preliminaries
@@
A homotopy $[0,1]\times M\ni(s,p)\mapsto H(s,p)$ is said to be an isotopy when $p\mapsto H(s,p)$ is a topological embedding for all $s\in[0,1]$. A homeomorphism $\psi:\mathbb{R}^{n}\to\mathbb{R}^{n}$ is said to be a stable homeomorphism when it is a finite composition of homeomorphisms that equal the identity map on some—not necessarily the same—non-empty open subset of $\mathbb{R}^{n}$. We denote by $\mathrm{Homeo}(\mathbb{R}^{n};\mathbb{R}^{n})$ the group (under composition) of homeomorphisms $\psi:\mathbb{R}^{n}\to\mathbb{R}^{n}$. Additionally, we denote by $\mathrm{Homeo}^{+}(\mathbb{R}^{n};\mathbb{R}^{n})\subset\mathrm{Homeo}(\mathbb{R}^{n};\mathbb{R}^{n})$ the subgroup of all orientation-preserving [^1] The reader that is unfamiliar with the notion of orientation is first directed to [ref:guillemin2010differential, Lee2] to see orientations in the smooth case. The topological definition relies on algebraic topology, see for instance [Hatcher, Sec. 3.3]. homeomorphisms.
A function $\alpha:\mathbb{R}_{\geq 0}\to\mathbb{R}_{\geq 0}$ is said to be of class $\mathcal{K}$ when $\alpha$ is strictly increasing and $\alpha(0)=0$. If, additionally $\lim_{s\to+\infty}\alpha(s)=+\infty$, then, $\alpha$ is of class $\mathcal{K}_{\infty}$. These type of functions are called comparison functions, e.g., see [hahn1967stability, ref:kellett2014compendium].
@@ 1.2   Dynamical systems
In this note we study deterministic, finite-dimensional, time-invariant, continuous, (global) semi-dynamical systems comprised of the triple $\Sigma:=({M}^{n},\varphi,\mathbb{R}_{\geq 0})$. Here, ${M}^{n}$ will be a smooth $n$-dimensional manifold diffeomorphic to $\mathbb{R}^{n}$, which we denote by ${M}^{n}\simeq_{d}\mathbb{R}^{n}$, and $\varphi:\mathbb{R}_{\geq 0}\times{M}^{n}\to{M}^{n}$ is a global semiflow, that is, a continuous map that satisfies for any $p\in{M}^{n}$: $(i)$ $\varphi(0,p)=p$ (the identity axiom); and $(ii)$ $(\varphi(s,\varphi(t,p))=\varphi(t+s,p)$ $\forall s,t\in\mathbb{R}_{\geq 0}:=\{t\in\mathbb{R}:t\geq 0\}$ (the semigroup axiom). We will usually write $\varphi^{t}(\cdot)$ instead of $\varphi(t,\cdot)$. In particular, we study semiflows generated by continuous vector fields over ${M}^{n}$, that is, when $\varphi$ satisfies
@@
where $X$ is a continuous section, that is, the map $X:M^{n}\to TM^{n}$ is continuous and satisfies $\pi\circ X=\mathrm{id}_{M^{n}}$ for $\pi$ the canonical projection $(p,v)\mapsto\pi(p,v)=p$. When $M^{n}=\mathbb{R}^{n}$, then, the identification $T\mathbb{R}^{n}\simeq\mathbb{R}^{n}\times\mathbb{R}^{n}$ results in being able to discuss vector fields as simply self-maps of $\mathbb{R}^{n}$. On $\mathbb{R}^{n}$, we will work a lot with the “canonical” differential equation $\dot{x}=-x$, but also with the identity map $x\mapsto x$, thus, to avoid notational confusion, we denote the corresponding canonical vector field by $-\partial_{x}$, and not $-\mathrm{id}_{\mathbb{R}^{n}}$.
The focus on semiflows instead of flows allows us to look at sufficiently regular discontinuous vector fields as well. This is relevant to control theory, as the introduction of feedback usually results in a closed-loop vector field that cannot be assumed to be continuous (e.g., think of optimal control [^2] One might also think of topological obstructions, however, as will become clear below, the type of discontinuities we consider do not allow for overcoming those obstructions, in general [ref:ryan1994brockett].). This choice of setting is also motivated by other recent work. For instance, the main research question in [ref:ozaslan2024exponential] is: “Given an exponentially stable optimization algorithm, can it be modified to obtain a finite/fixed-time stable algorithm?”. They provide some sufficient conditions, we will show that a more generic, yet less constructive viewpoint is also possible.
@@
where the set-valued map $F:\mathbb{R}^{n}\to 2^{\mathbb{R}^{n}}$, with $2^{X}$ denoting the power set of $X$, is in some precise sense related to $X$. The intuition is to pass from a irregular single-valued map, to a more regular set-valued map that contains the original behaviour.
Let $\lambda^{d}$ denote the Lebesgue measure on $\mathbb{R}^{d}$, then, solutions to these differential inclusions, are absolutely continuous curves $\xi\in\mathrm{AC}(\mathcal{I}\subseteq\mathbb{R};\mathbb{R}^{n})$ such that $\dot{\xi}(t)\in F(\xi(t))$ for $\lambda^{1}$-a.e. $t\in\mathcal{I}$. Typically, $F$ is assumed to be upper semi-continuous and compact, convex valued. With those assumptions in mind, then, under mild conditions on $X$, a valuable solution framework follows by applying Filippov’s operator $\mathcal{F}$, that is,
$x\mapsto\mathcal{F}[X](x):=\bigcap_{\delta>0}\bigcap_{N\in\{A\subset\mathbb{R}^{n}:\lambda^{n}(A)=0\}}\overline{\mathrm{conv}}\,X\left(\{x\in\mathbb{R}^{n}:\|x\|_{2}<\delta\}\setminus N\right).$
Then, solutions to
$\dot{x}\in\mathcal{F}[X](x)$ (1.3)
are understood to be “generalized” solutions to $\dot{x}=X(x)$, usually called Filippov solutions, i.e., solutions to (1.3) are absolutely continuous curves $\xi:\mathcal{I}\to\mathbb{R}^{n}$ such that $\dot{\xi}(t)\in\mathcal{F}[X](\xi(t))$ for $\lambda^{1}$-a.e. $t\in\mathcal{I}$.
For more on differential inclusions and discontinuous dynamical systems, we point the reader to [ref:filippov1988differential], [ref:bacciotti2005liapunov, Ch. 1] and [ref:cortes2008discontinuous].
@@ 1.3   Stability
To characterize stability under a differential inclusion (1.2), we need a few concepts. Starting with the regular case, for simplicity, let the vector field $F$ be single-valued and smooth. In that case, $F$ generates a flow, denoted $\varphi(\cdot;F)$. A point $x^{\star}\in\mathbb{R}^{n}$ is an equilibrium point of $F$ when $F(x^{\star})=0$ or equivalently $\varphi^{t}(0;F)=0$ $\forall t\in\mathbb{R}$. We will set $x^{\star}$ to be 00, unless stated otherwise. Then, 00 is said to be globally asymptotically stable (GAS) (under $F$) when
@@
Although we work with $\mathbb{R}^{n}$, or $M^{n}\simeq_{d}\mathbb{R}^{n}$, we take a topological approach akin to [ref:bhatiahajek2006local] and hence the definition of GAS is as above. When working with metrics, one typically says more, e.g., see [ref:wilson1969smoothing, ref:lin1996smoothV2], there, a metric is used to characterize how Lyapunov stability can be truly turned non-local (e.g., let $B_{r}(x)$ be some $r$-metric ball at $x$, then, there is a some $\mathcal{K}_{\infty}$ function $\delta$ such that for any $\varepsilon$ we have that $\varphi^{t}(B_{\delta(\varepsilon)}(x))\subseteq\varphi^{t}(B_{\varepsilon}(x))$ $\forall t\geq 0$.) In general these definitions are not equivalent, however, for being GAS on $\mathbb{R}^{n}$, they are [ref:andriano1997global]. $\circ$
Due to the work of Lyapunov [ref:liapunov1892general], we know that to reason about stability, it is worthwhile to look for “potential functions” that capture stability, illustrated by the fact that his theory effectively replaced the definitions of stability. Specially, we look for a function $V\in C^{\infty}(\mathbb{R}^{n};\mathbb{R}_{\geq 0})$, satisfying the following properties:
@@
    $\langle\nabla V(x),F(x)\rangle<0$ for all $x\in\mathbb{R}^{n}\setminus\{0\}$;
@@
    $V$ is radially unbounded, that is, $V(x)\to+\infty$ for $\|x\|_{2}\to+\infty$.
[…]
@@
A pair of functions $(V,W)\in C^{0}(\mathbb{R}^{n};\mathbb{R}_{\geq 0})$, with $V$ being $C^{\infty}$-smooth on $\mathbb{R}^{n}$ and $W$ being $C^{\infty}$-smooth on $\mathbb{R}^{n}\setminus\{0\}$, is said to be a $C^{\infty}$-smooth strong Lyapunov pair for the vector field $F$ as in (1.2), provided that $F$ is upper semi-continuous and compact, convex valued, plus, the following conditions hold:
@@
The preposition “strong” comes from the fact that we look at all solutions satisfying (1.2). Similarly, one could require that at least one solution is stable and construct a “weak” version of Definition 1.1.
For further references, and generalizations of Lyapunov’s stability theory, we point the reader to [ref:bhatia1970stability, ref:sontag2013mathematical, ref:bacciotti2005liapunov, ref:bhatiahajek2006local, ref:goebel2012hybrid].
Before closing this section, we explicitly illustrate why working with $C^{\infty}$ or even $C^{0}$ vector fields is arguably overly restrictive when interested in qualitative stability questions. We start with a simple example.
@@
$\dot{x}=X_{1}(x):=-\mathrm{sgn}(x):=\begin{cases}1\quad&x<0\\ 0\quad&x=0\\ -1\quad&x>0\end{cases}.$ (1.4)
Now, consider the map $\varphi_{1}:\mathbb{R}_{\geq 0}\times\mathbb{R}\to\mathbb{R}$ defined through
@@
One can check that $\varphi$ is a global semiflow, describing a solution (e.g., in the sense of Carathéodory, Krasovskii or Filippov [ref:filippov1988differential]) to (1.4)—a discontinuous vector field. In particular, note that $t\mapsto\varphi_{1}^{t}(x)\in\mathrm{AC}$ since $\varphi_{1}^{t}(x)=\varphi_{1}^{0}(x)+\int^{t}_{0}-\mathrm{sgn}(\varphi_{1}^{\tau}(x))\mathrm{d}\tau$ for any $t\geq 0$. Regarding stability, consider the $C^{\infty}$ Lyapunov function $x\mapsto V_{1}(x):=\tfrac{1}{2}x^{2}$ and find that $\nabla V_{1}(x)X_{1}(x)=-|x|<0$ on $\mathbb{R}\setminus\{0\}$ (consider a Huber loss to find a valid function $W_{1}$ in the sense of Definition 1.1). This already shows that the existence of a smooth Lyapunov function, asserting that the origin is GAS, does not imply the existence of a flow, nor does it imply that convergence to 00 is merely asymptotic. Now let $X_{2}(x):=-\partial_{x}$ and define $[0,1]\ni s\mapsto X(\cdot;s):=(1-s)X_{1}+sX_{2}$. Then, for any $s\in[0,1]$ we have that $\nabla V(x)X(x;s)<0$ on $\mathbb{R}\setminus\{0\}$. We know that the flow corresponding to $X_{2}$ is $(t,x)\mapsto\varphi_{2}^{t}(x):=e^{-t}x$. Better yet, by direct integration, we find that a global semiflow corresponding to $X(\cdot;s)$ becomes
@@
Indeed, $s\mapsto\varphi(\cdot;s)$ parametrizes a homotopy, along global semiflows such that 00 is GAS, from $\varphi_{1}$ to $\varphi_{2}$. Regarding $\lim_{s\to 0^{+}}\varphi(\cdot;s)$, use that $e^{-st}=1-st+\sum^{\infty}_{n=2}((-st)^{n}/n!)$ to recover (1.5). It is interesting to note that although (1.5) is the unique (AC) solution to $\dot{x}\in\mathcal{F}[X_{1}](x)$, the solution to $\dot{x}\in\mathcal{F}[-X_{1}](x)$ is not unique, stability is key. $\circ$
Next, we provide an example that will appear later.

### v2

@@ Abstract
Keywords—dynamical systems, homotopy invariant, stability
@@ 1   Introduction
In the context of what we call today Conley index theory (see Section A), Conley posed the following “converse question” in the 1970s: “To what extent does the homotopy index [Conley index] itself determine the equivalence class of isolated invariant sets which are related by continuation?” [ref:conley1978isolated, p. 83]. Recently, Kvalheim proved that uniquely integrable $C^{0}$ vector fields, on a $C^{\infty}$ manifold $M$, rendering a compact set $A\subseteq M$ asymptotically stable, are homotopic on an open neighbourhood $U\supseteq A$ such that throughout the homotopy the vector fields do not vanish on $U\setminus A$ [ref:kvalheim2022obstructions, Thm. 1]. Connecting this result to Conley, a follow-up question (revitalized) by Kvalheim—which is the central question of this note—is the following.
Question:Are dynamical systems that render a set $A$ asymptotically stable, homotopic through dynamical systems that preserve this notion of stability?
This question, in one form or another, inspired several works, for instance, [ref:reineck1992continuation, ref:mrozek2000conley, ref:JongeneelSchwan2024TAC]. Here, we will elaborate on commentary by the author in [ref:JongeneelSchwan2024TAC]. Specifically, in the seminal paper “Asymptotic stability equals exponential stability, and ISS equals finite energy gain—if you twist your eyes” from 1999, Grüne, Sontag and Wirth showed that asymptotic stability “equals” exponential stability in the sense that if an equilibrium point is asymptotically stable under some vector field on $\mathbb{R}^{n}$, then, there is a suitable change of coordinates rendering this point exponentially stable [ref:grune1999asymptotic]. Such a change of coordinates is understood to be instantaneous. However, by leveraging their work, we show that asymptotic stability can be continuously “transformed” into exponential stability, while preserving asymptotic stability throughout the transformation, see Theorem 3.6. Differently put, asymptotic stability equals exponential stability—not only if you twist your eyes, but while you twist your eyes. The key to this is to observe that we can select the transformation from [ref:grune1999asymptotic] to be an orientation-preserving homeomorphism on $\mathbb{R}^{n}$, not just any homeomorphism cf[ref:JongeneelSchwan2024TAC, Sec. III], see the proof of Proposition 3.2.
This result provides a partial solution to Conley’s converse question as it turns out that the asymptotically stable systems under consideration can be continuously transformed into the same exponentially stable system and hence, by transitivity, into each other. Concurrently, we discuss extensions to discontinuous vector fields throughout, plus we illustrate how to go about extensions to ISS (see Section 4). We also discuss intimate connections with optimization and optimal transport (e.g., see Example 3.4 and 5.1).
@@ 1.1   Related work
It can be argued that questions of the form above emerged from studies aimed at classifying manifolds, maps and so forth. A successful, yet coarse, resolution has been found in the study of these objects up to homotopy, e.g., motivated early on by the fundamental group being homotopy invariant, Hopf’s degree theorem, CW complexes, intractability of topological equivalence and more at the intersection of topology and dynamical systems.
@@
We cannot do justice here to the wealth of work in this area, but inspired by Andronov, Pontryagin, Thom, Peixoto, Birkhoff, Milnor and several others, it was in particular Smale highlighting that this intersection is fruitful for both fields, see [ref:hirsch2012topology] for context. Indeed, after seminal work by Smale on the qualitative classification of dynamical systems [ref:smale1967differentiable], homotopy results concerned with Morse-Smale vector fields appeared, e.g., see [ref:asimov1975homotopy, ref:newhouse1976, ref:franks1979morse]. Similarly, one can study if gradient vector fields are homotopic through gradient vector fields [ref:parusinski1990gradient], see also [ref:reineck1991continuation, ref:kvalheim2022obstructions] for work close in spirit to ours.
@@
In control theory, these tools (i.e., certain homotopy invariants) were used to construct necessary conditions for feedback controllers to exist, i.e., if a desirable (closed-loop) dynamical system belongs to a certain homotopy class, then, there must be at least some feedback that renders the control system a member of this class, e.g., see [ref:brockett1983asymptotic, ref:krasnosel1984geometrical, ref:zabczyk1989, ref:Coron1990].
@@
These results have in common that the original objects are “homotoped” to something simple and “canonical” where we do our analysis. For instance, in the context of control systems, when our goal is stabilization of the origin on $\mathbb{R}^{n}$ through feedback, then, the canonical differential equation is $\dot{x}=-x$ and we would like our closed-loop dynamical system to be in some qualitative sense equivalent to this equation. This is precisely how the “index condition” by Krasnosel’skiĭ and Zabreĭko is derived [ref:krasnosel1984geometrical, Sec. 52]. Clearly, necessary conditions of this form are only as valuable as the homotopy class is distinctive, e.g., although the index condition is powerful, it cannot differentiate between $\dot{x}=x$ and $\dot{x}=-x$ on $\mathbb{R}^{2}$. This is precisely what motivates us, we expect that understanding Conley’s converse question can lead to stronger necessary conditions for stabilization through continuous feedback.
@@
Besides possibly stronger necessary conditions for continuous stabilization, the study of Conley’s converse question directly relates to understanding topological properties of spaces of stable dynamical systems (as was motivation of prior work [ref:JongeneelSchwan2024TAC]). This is of practical importance as several frameworks in optimal control and reinforcement learning aim to optimize over precisely such a space, e.g., moving from a pre-stabilized system to an optimally controlled system, we point to [ref:Fazel_18, ref:furieri2024learning] and references therein.
@@
Before we continue, we recall that Conley’s converse question is not trivial.
@@
Both $A(0)$ and $A(1)$ correspond to global asymptotically stable systems, yet, for $s=\tfrac{1}{2}$, the system $\dot{x}=A(s)x$ is unstable. Hence, we cannot just construct straight-line homotopies between stable vector fields and expect that stability is preserved. Instead, we know from [ref:JongeneelSchwan2024TAC] that for vector fields with convex Lyapunov functions we should homotope via the canonical ODE $\dot{x}=-x$. Explicitly, consider the following path of vector fields defined by
$H(x;s):=\begin{pmatrix}-1&\max\{0,(1-2s)10\}\\ \max\{0,(2s-1)10\}&-1\end{pmatrix}x,\quad s\in[0,1].\penalty 9999\hbox{}\nobreak\hfill\quad\hbox{$\circ$}$
Notation
@@ 1.2   Dynamical systems
We study semi-dynamical systems comprised of the triple $\Sigma:=({M}^{n},\varphi,\mathbb{R}_{\geq 0})$. Here, ${M}^{n}$ will be a smooth $n$-dimensional manifold (frequently, diffeomorphic to $\mathbb{R}^{n}$, which we denote by ${M}^{n}\simeq_{d}\mathbb{R}^{n}$,) and $\varphi:\mathbb{R}_{\geq 0}\times{M}^{n}\to{M}^{n}$ is a (global) semiflow, that is, a continuous map that satisfies for any $p\in{M}^{n}$: $(i)$ $\varphi(0,p)=p$; and $(ii)$ $(\varphi(s,\varphi(t,p))=\varphi(t+s,p)$ $\forall s,t\in\mathbb{R}_{\geq 0}:=\{t\in\mathbb{R}:t\geq 0\}$. We will usually write $\varphi^{t}(\cdot)$ instead of $\varphi(t,\cdot)$. In particular, we study (forward complete) semiflows generated by vector fields over ${M}^{n}$, that is, when $\varphi$ satisfies
@@
where $X$ is a continuous section, that is, the map $X:M^{n}\to TM^{n}$ is continuous and satisfies $\pi\circ X=\mathrm{id}_{M^{n}}$ for $\pi$ the projection $(p,v)\mapsto\pi(p,v)=p$. When $M^{n}=\mathbb{R}^{n}$, the identification $T\mathbb{R}^{n}\simeq_{d}\mathbb{R}^{n}\times\mathbb{R}^{n}$ allows for discussing vector fields as self-maps of $\mathbb{R}^{n}$. On $\mathbb{R}^{n}$, we will frequently work with the “canonical” ODE $\dot{x}=-x$, but also with the identity map $x\mapsto x$. To avoid confusion, we denote the canonical vector field by $-\partial_{x}$, and not $-\mathrm{id}_{\mathbb{R}^{n}}$.
The focus on semiflows instead of flows allows us to look at sufficiently regular discontinuous vector fields. This is relevant, as the introduction of feedback usually results in a closed-loop vector field that cannot be assumed to be continuous (e.g., think of optimal control [^2] One might also think of topological obstructions, however, the type of discontinuities we consider do not allow for overcoming those obstructions, in general [ref:ryan1994brockett].). This choice of setting is also motivated by other recent work. For instance, the research question in [ref:ozaslan2024exponential] is: “Given an exponentially stable optimization algorithm, can it be modified to obtain a finite/fixed-time stable algorithm?”. Some sufficient conditions are provided in [ref:ozaslan2024exponential], we will show a generic, yet less constructive viewpoint.
@@
where the set-valued map $F:\mathbb{R}^{n}\to 2^{\mathbb{R}^{n}}$, with $2^{S}$ denoting the power set of a set $S$, is in some precise sense related to $X$. The rationale is to pass from a irregular single-valued map, to a more regular set-valued map.
Let $\lambda^{d}$ denote the Lebesgue measure on $\mathbb{R}^{d}$, then, solutions $\mathcal{I}\ni t\mapsto\xi(t)$ to these differential inclusions, are absolutely continuous (AC) on each compact subinterval of $\mathcal{I}$ and such that $\dot{\xi}(t)\in F(\xi(t))$ for $\lambda^{1}$-a.e. $t\in\mathcal{I}$. Typically, $F$ is assumed to be upper semi-continuous and compact, convex valued. With those assumptions in mind, then, under mild conditions on $X$, a valuable solution framework follows by applying Filippov’s operator $\mathcal{F}$, that is,
$x\mapsto\mathcal{F}[X](x):=\bigcap_{\delta>0}\bigcap_{N\subset\mathbb{R}^{n}:\lambda^{n}(N)=0}\overline{\mathrm{co}}\,X\left(B^{n}(x,\delta)\setminus N\right),$
for $\overline{\mathrm{co}}(\cdot)$ the closure of the convex hull and $B^{n}(x,\delta):=\{z\in\mathbb{R}^{n}:\|x-z\|_{2}<\delta\}$. Then, solutions to $\dot{x}\in\mathcal{F}[X](x)$ are understood to be “generalized” solutions to $\dot{x}=X(x)$, called Filippov solutions, and such that $\dot{\xi}(t)\in\mathcal{F}[X](\xi(t))$ for $\lambda^{1}$-a.e. $t\in\mathrm{dom}(\xi)$. For references, see [ref:filippov1988differential], [ref:bacciotti2005liapunov, Ch. 1] and [ref:cortes2008discontinuous].
@@ 1.3   Stability
In this section we will characterize stability under (1.2). Starting with the regular case, for simplicity, let the vector field $F$ be single-valued and smooth. In that case, $F$ generates a flow, denoted $\varphi(\cdot;F)$. A point $x^{\star}\in\mathbb{R}^{n}$ is an equilibrium point of $F$ when $F(x^{\star})=0$ or equivalently $\varphi^{t}(0;F)=0$ $\forall t\in\mathbb{R}$. We will set $x^{\star}:=0$, unless stated otherwise. Then, 00 is said to be globally asymptotically stable (GAS) (under $F$) when
@@
Although we merely work with $\mathbb{R}^{n}$ (or $M^{n}\simeq_{d}\mathbb{R}^{n}$) we take a topological approach akin to [ref:bhatiahajek2006local] and hence GAS is defined as above. One can turn Property (s.i) truly global, however, e.g., see [ref:wilson1969smoothing, ref:lin1996smoothV2]. Specifically, let $B_{r}(x)$ be some $r$-metric ball centered at $x$, then, require the existence of $\delta\in\mathcal{K}_{\infty}$ such that for any $\varepsilon\geq 0$ we have $\varphi^{t}(B_{\delta(\varepsilon)}(x))\subseteq B_{\varepsilon}(x)$ $\forall t\in\mathbb{R}_{\geq 0}$. In general these definitions are not equivalent, however, for 00 being GAS on $\mathbb{R}^{n}$, they are [ref:andriano1997global]. $\circ$
Due to the work of Lyapunov [ref:liapunov1892general], we know that to reason about stability, it is worthwhile to look for “potential functions” that capture stability, illustrated by the fact that his theory effectively replaced the definitions of stability. Specifically, we look for $V\in C^{\infty}(\mathbb{R}^{n};\mathbb{R}_{\geq 0})$ satisfying:
@@
    $V$ is radially unbounded, that is, $V(x)\to+\infty$ for $\|x\|\to+\infty$; and
@@
    $\langle\nabla V(x),F(x)\rangle<0$ for all $x\in\mathbb{R}^{n}\setminus\{0\}$.
Property (V.ii) implies sublevel set compactness. We call such a function a (smooth, strict and proper) Lyapunov function (with respect to the pair $(F,0)$). This work is about GAS, so we will omit “strict” and “proper” from now on. Then, based on converse theory, e.g., see [ref:Massera1956, ref:kurzweil1963inversion, ref:wilson1969smoothing, ref:fathi2019smoothing], we can appeal to the celebrated theorem stating that 00 is GAS if and only if there is a (corresponding) smooth Lyapunov function [ref:bacciotti2005liapunov, Thm. 2.4].
@@
A pair of continuous functions $(V,W)$, with $V\in C^{\infty}(\mathbb{R}^{n};\mathbb{R}_{\geq 0})$ and $W\in C^{\infty}(\mathbb{R}^{n}\setminus\{0\};\mathbb{R}_{\geq 0})$, is said to be a $C^{\infty}$ strong Lyapunov pair for $F$, as in (1.2), provided that $F$ is upper semi-continuous and compact, convex valued, plus, the following hold:
@@
The preposition “strong” comes from the fact that we look at all solutions satisfying (1.2). For further references, and generalizations of Lyapunov’s stability theory, we point the reader to [ref:bhatia1970stability, ref:sontag2013mathematical, ref:bacciotti2005liapunov, ref:bhatiahajek2006local, ref:goebel2012hybrid].
Before closing this section, we explicitly illustrate why working with $C^{\infty}$ or even $C^{0}$ vector fields is arguably overly restrictive when interested in qualitative stability questions. We provide two examples that will return.
@@
$\dot{x}=X_{1}(x):=-\mathrm{sgn}(x),$ (1.3)
with $\mathrm{sgn}(0):=0$. Now, consider the map $\varphi_{1}:\mathbb{R}_{\geq 0}\times\mathbb{R}\to\mathbb{R}$ defined through
@@
One can check that $\varphi_{1}$ is a semiflow, describing a solution (e.g., in the sense of Filippov) to (1.3). Regarding stability, consider the $C^{\infty}$ Lyapunov function $x\mapsto V_{1}(x):=\tfrac{1}{2}x^{2}$ and see that $\nabla V_{1}(x)X_{1}(x)=-|x|<0$ on $\mathbb{R}\setminus\{0\}$ (consider a Huber loss to find a valid $W_{1}$ in the sense of Definition 1.3). This already shows that the existence of a smooth Lyapunov function, asserting that the origin is GAS, does not imply the existence of a flow, nor does it imply that convergence to 00 is merely asymptotic. Now let $X_{2}(x):=-\partial_{x}$, define $[0,1]\ni s\mapsto X(\cdot;s):=(1-s)X_{1}+sX_{2}$ and see that for and $s\in[0,1]$ we have $\nabla V(x)X(x;s)<0$ on $\mathbb{R}\setminus\{0\}$. The semiflow corresponding to $X(\cdot;s)$ becomes
@@
Then, $s\mapsto\varphi(\cdot;s)$ parametrizes a homotopy, along semiflows such that 00 remains GAS. For $\lim_{s\to 0^{+}}\varphi(\cdot;s)$, use $e^{-st}=1-st+\sum^{\infty}_{n=2}((-st)^{n}/n!)$. $\circ$